# FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models

### v3

@@ Abstract
In this work, we propose using a unified representation, termed Factorized Features, for low-level vision tasks, where we test on Single Image Super-Resolution (SISR) and Image Compression. Motivated by the shared principles between these tasks, they require recovering and preserving fine image details—whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition as well as an explicit formulation of frequencies to capture structural components and multi-scale visual features in images, which addresses the core challenges of both tasks. We replace the representation of prior models from simple feature maps with Factorized Features to validate the potential for broad generalizability. In addition, we further optimize the pipelines by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression and super-resolution. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA. Please see our project page at jayisaking.github.io/FIPER.
@@ 1 Introduction
Figure 1: The Proposed Factorized Features. With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression. Note that in this figure $\alpha=\{1,16\}$ and $\psi=\text{sin}(\cdot)$.
@@
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches [52, 23] laid the foundation for SISR, which was later enhanced with GAN-based methods [40] for improved perceptual realism. Follow-up Transformer-based networks [14] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [21, 17, 18], which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR? On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [35, 82, 79] lay the groundwork. The emerging learned image compression models [8, 3, 30, 55, 67, 19], compression algorithms of which mostly follow the pixel-space transform coding  [8, 28] paradigm, then introduce a variety of networks  [83, 55] or difference losses [31] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
@@
Based on the aforementioned analysis, it becomes clear that although these tasks appear to be distinct, they share mutual similarities in two key aspects: (1) The tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) They aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [11, 68, 9, 26, 7, 12, 27], we propose a unified representation, Factorized Features, with generalizable Coefficient Backbone and Basis Transformer for learned coefficient and basis, respectively. This approach explicitly captures multi-scale visual features and repetitive structral components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation.
@@
Finally, we propose merging the bases for multi-image processing by consolidating multiple bases into one through a network, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
@@
  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in Sec. 3.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
  • We generalize such image descriptor to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
@@ 2 Related Works
Super-Resolution (SR). Image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, which is crucial for computer vision. Early CNN-based methods [23, 52, 95, 76, 36] introduced residual [41, 54, 81, 94, 93, 20, 51] and recursive learning [18, 77]. GAN-based approaches [40, 41, 84, 85] improved perceptual quality but faced spatial locality constraints, leading to Transformer-based SISR models [14, 46]. SwinIR [48] integrated window attention with Swin Transformer [58], inspiring further advancements [21, 96, 91, 20]. Hybrid models like CRAFT [43], DAT [17], and HAT [15] optimize feature aggregation, while RGT [18] enhances spatial details efficiently.
Image Compression (IC). Deep learning surpasses traditional compression methods like JPEG [82] and JPEG-2000 [79], with CNN-based [8, 4, 29], Transformer-based [38, 97, 86, 24], and GAN-based [61, 62, 25, 73] approaches improving performance. ELIC [32] introduced adaptive coding, LIC TCM [55] combined CNNs and Transformers, and eContextformer [39] leveraged spatial-channel attention. GroupedMixer [44] introduced token mixers, while Wavelet Conditional Diffusion [75] balanced perceptual quality and distortion.
@@
3 Methods
In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [57], and Factor Fields [12], and delve into the motivation and how we resolve the weaknesses of these formulations by our Factorized Features in Sec. 3.1. Next, we discuss how to integrate such strategy into learned settings of Super-Resolution in Sec. 3.2 and Image Compression Sec. 3.3. Note that the background of learned image compression is provided in supplementary materials.
Figure 3: The overall pipeline of image super-resolution with our Factorized Features. Given a low-resolution image $I_{LR}$, we first extract the coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficients and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, and $W$ denote the scale factor, height, and width, respectively.
3.1 Formulation of Factorized Features
Fourier Transform has been long used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions. In practice, it often takes the form of Discrete Cosine Transform for real-valued 2D discrete signal $f$ with height $H$ and width $W$:
$f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right]$ (1)
@@ 3.1 Formulation of Factorized Features
$f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right]$ (2)
where $x<T$, $y<T$, and $\xi_{u,v}$ is phase shift. In image processing tasks, the goal is to learn $S$ and $\xi$, then use the formula above to reconstruct high-quality images from degraded inputs. To achieve this, [57] treats $S$ and $\xi$ as learnable parameters, allowing the model to optimize them for better reconstruction. Furthermore, the frequencies $u$ and $v$ are also designed to be promptable by the model, enabling dynamic adaptation to spatial-varying high- and low-frequency components.
@@
$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),$ (3)
where $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ and $N$ denotes the number of frequency components.
Spatially Variant Coefficient. Intuitively, the local frequency spectrum of a signal is spatially variant, meaning that different regions exhibit distinct spectral characteristics, e.g., certain regions of an image do not require high-frequency components. Therefore, in this work, we argue that coefficient $c_{i}$ should be spatial-varying, which is different from previous works, i.e.,
$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
Learned Non-uniform Basis. Previous works learn basis through frequencies ${u,v}$ and phase shift $\xi$[57, 47, 42]; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details; in other words, since low-frequency components typically require more energy [45], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details, and the predefined nature of these sinusoidal functions further restricts local flexibility, making fine-grained adjustments more challenging. As shown in Fig. 5, higher frequency components converge slower than their counterpart. Hence, in this work, we generate the entire basis map $b_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map and combine them together. In this work, the composition process happens in the end with a projection function $\mathcal{P}$ to form RGB images $\hat{I}$:
$\hat{I}(\mathbf{x})=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl{\}}\Bigl{)}.$ (5)
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Eq. 4 (b) Eq. 6 (c) Eq. 7 with only $\psi$ (d) Eq. 7 with both $\psi$ and $\alpha$
Figure 4: Visualization of different factorization methods.
Coordinate Transformation. The success of Fourier-based representation originates from its sinusoidal decomposition. By using such periodic functions, one can effectively analyze and reconstruct signals with the enforcement of high- and low-frequency components. However, the formulation in Eq. 5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in Fig. 4 we witness performance loss due to lack of such guidance. Inspired by position encoding [65, 11, 68, 7, 9] and Fourier feature mapping [78], we introduce coordinate transformation function $\gamma_{i}$ on $\mathbf{x}$ before sampling basis $b_{i}$:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, visualized in Fig. 2.a. We argue that the function explicitly imposes patch-like periodic constraints and, therefore, enforces models to learn such repetitive patterns.
Multi-frequency Modulation. [72] has shown that neural networks are biased towards learning low-frequency content. Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in Fig. 4:
@@
Alternatively, Eq. 7 can be interpreted as making the spatial coordinates $x,y$ in Eq. 1 learnable instead of the frequency components $u,v$. To validate this, we conduct experiments (Fig. 5) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
To this end, our full Factorized Features representation is formulated in Eq. 7.
Figure 5: In this figure, we validate the claim that fixed sinusoidal functions attend more to low-frequency parts. Learnable basis denotes the Learned Non-uniform basis in Sec. 3.1. And we can see that Learnable Coordinate $x$ has the same effect, verifying our conclusion.
Figure 6: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec. 3.2 as information-recovery prior, detailed in Sec. 3.3 (c) Expanding on (b),  Sec. 3.3.1 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.

### v4

# FIPER: Factorized Features for Robust Image Super-Resolution and Compression

@@ Abstract
In this work, we propose using a unified representation, termed Factorized Features, for low-level vision tasks, where we test on Single Image Super-Resolution (SISR) and Image Compression. Motivated by the shared principles between these tasks, they require recovering and preserving fine image details, whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression. Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition as well as an explicit formulation of frequencies to capture structural components and multi-scale visual features in images, which addresses the core challenges of both tasks. We replace the representation of prior models from simple feature maps with Factorized Features to validate the potential for broad generalizability. In addition, we further optimize the compression pipeline by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression. Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA. Project page: https://jayisaking.github.io/FIPER/
@@ 1 Introduction
Figure 1: In this work, we propose to represent images by Factorized Features, contrasting with prior methods that either (a) rely on fixed-basis decomposition[66, 39] or (b) introduce various complex model blocks [15, 71] such as sparse attention, dense skip connection, etc. As highlighted in the orange boxes, (a) successfully captures periodic line structures but suffers from lower fidelity, while (b) cannot recover those repeating patterns without explicit frequency modeling. In contrast, our method (c) uses learned decomposition with generalizable bases to deliver high-fidelity reconstruction while preserving periodic structures. We also present LAM [32] and Diffusion Index (DI) [32] for reconstructing the red-boxed patch, with its pixel-importance map shown on the middle row; a higher DI indicates more pixel utilization.
Single-image super-resolution (SISR) aims to recover high-quality images from low-resolution inputs, with accuracy depending on precise restoration of fine details and geometric correspondences (e.g., stripes, grids, textures). CNN pioneers [24, 59] were followed by GAN-based methods for perceptual realism [47] and Transformer networks for long-range context [14]. Swin-based variants further boosted performance [22, 18, 19, 71], inspiring ever-heavier architectures. These prior works optimize network design while neglecting the underlying image-content representations themselves. On the other hand, decomposition-based methods[66, 44, 39] seek to model the frequency components directly through adding Fourier- or Wavelet-like bottlenecks in the middle. Although these designs can indeed better capture the recurring visual patterns in learning effective SR representations than aforementationed methods, as illustrated in fig.˜1, they still suffer from poor perceptual quality.
This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
Conventional multi-scale pyramids aggregate features at several resolutions, but they store all activations densely, do not disentangle frequency content, and cannot be shared directly across tasks. Instead, our factorized descriptor decomposes images into generalizable basis $\times$ coefficient pairs, providing an explicit multi-frequency handle through different modulations.
@@
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [41, 95, 92] lay the groundwork. The emerging learned image compression models [8, 3, 34, 62, 80, 20], compression algorithms of which mostly follow the pixel-space transform coding  [8, 31] paradigm, then introduce a variety of networks  [96, 62, 106, 67, 27] or difference losses [35] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
@@
Based on the aforementioned analysis, it becomes clear that although these tasks appear to be distinct, they share mutual similarities in two key aspects: (1) The tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) They aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity. Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [11, 81, 9, 28, 7, 12, 30], we propose a unified representation, Factorized Features, with generalizable Coefficient Backbone and Basis Transformer for learned coefficient and basis, respectively. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation.
@@
Finally, we also propose consolidating multiple bases into one through a network for multi-image processing, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
@@
  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in section˜4.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
  • We generalize such image descriptors to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
@@ 2 Related Works
Super-Resolution (SR). Image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, which is crucial for computer vision. Early CNN-based methods [24, 59, 113, 89, 43] introduced residual [48, 61, 94, 112, 111, 21, 58] and recursive learning [19, 90]. GAN-based approaches [47, 48, 97, 98] improved perceptual quality but faced spatial-locality constraints, leading to Transformer-based SISR models [14, 53]. SwinIR [55] integrated window attention with Swin Transformer [69], inspiring further advancements [22, 114, 109, 21, 71]. Hybrid models such as CRAFT [50], DAT [18], and HAT [15] optimize feature aggregation, while RGT [19] enhances spatial details efficiently. Recent efficient Transformer variants such as Restormer [105] and Uformer [99] demonstrate that lightweight attention or activation-free blocks can match or surpass heavier designs. Diffusion-based methods such as DDRM [42] provide an orthogonal line of plug-and-play restoration priors, with recent advances extending to video restoration [104], reference-based face restoration [37], and inpainting tasks [64]. Additionally, learned pipeline approaches [68] reverse the image formation process for reconstruction tasks.
Spectral Coordinate Transforms. A separate line of work seeks to encode coordinates rather than deepen the backbone. Fourier Features [91, 78] map input positions to a fixed harmonic basis, enabling MLPs to model high-frequency detail. CoordConv [65] expands convolutional layers with raw $(x,y)$ channels to facilitate spatial reasoning. On the other hand, sawtooth coordinate transformation $\gamma(\cdot)$, of our Factorized Features, folds space into a piece-wise linear triangle wave and produces evenly spaced spectral peaks (see section˜4.1) that favor sparse and low-rank Factorized Features. Unlike the above encoding, our basis-coefficient decomposition paired with $\gamma$ can be trained end-to-end and reused directly in both super-resolution and compression pipelines.
Image Compression (IC). Deep learning surpasses traditional codecs such as JPEG [95] and JPEG-2000 [92], with CNN-based [8, 4, 33], Transformer-based [45, 115, 100, 25, 27, 106, 67, 29], and GAN-based [73, 74, 26, 86] approaches improving performance. ELIC [36] introduced adaptive coding; LIC TCM [62] combines CNNs and Transformers; eContextformer [46] leverages spatial–channel attention. GroupedMixer [51] introduces token mixers, while Wavelet Conditional Diffusion [88] balances perceptual quality and distortion. Generative codecs such as HiFiC [77] pursue high-fidelity reconstructions at extremely low bitrates and offer a complementary direction to likelihood-based models.
@@
3 Frequency Decomposition Preliminary
In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [66], and Factor Fields [12], and then generalize them to a concise formulation of basis-coefficient decomposition as in eq.˜3.
The Fourier Transform has long been used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions. In practice, it often takes the form of the Discrete Cosine Transform for a real-valued 2D discrete signal $f$ with height $H$ and width $W$:
$\small f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right],$ (1)
@@ 3 Frequency Decomposition Preliminary
$\small f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right],$ (2)
where $x<T$, $y<T$, and $\xi_{u,v}$ is phase shift.
In image processing tasks, the goal is to learn $S$ and $\xi$, then use the formula above to reconstruct high-quality images from degraded inputs. To achieve this, [66] treats $S$ and $\xi$ as learnable parameters, allowing the model to optimize them for better reconstruction. Furthermore, the frequencies $u$ and $v$ are also designed to be promptable by the model, enabling dynamic adaptation to spatial-varying high- and low-frequency components.
@@
$\small f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),\ b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right],$ (3)
where $N$ denotes the number of frequency components.
In the next section, starting from the concise formulation in eq.˜3, we introduce our Factorized Features framework and elaborate on its spatially variant coefficients and adaptive basis functions, while showing its application to super-resolution and image compression.
4 Methods
In this section, we dive into the motivation and how we resolve the weaknesses of previous decomposition-based methods by our Factorized Features in section˜4.1. Next, we discuss how to integrate such a strategy into learned settings of Super-Resolution in section˜4.2 and Image Compression section˜4.3. We provide the background of learned image compression in the supplementary materials.
4.1 Formulation of Factorized Features
In this subsection, we focus on the analysis of the properties from different factorization components, and derive our final Factorized Features through these insights, starting from the simplest expression eq.˜3.
Spatially Variant Coefficient. Intuitively, the local frequency spectrum of a signal is spatially variant, meaning that different regions exhibit distinct spectral characteristics, e.g., certain regions of an image do not require high-frequency components. Therefore, in this work, we set coefficient $c_{i}$ to be spatial-varying, following [12], i.e.,
$\small f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
Learned Non-uniform Basis. Previous works[66, 54, 49] learn basis through frequencies u,v{u,v} and phase shift $\xi$; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details. The predefined nature of these sinusoidal functions restricts local flexibility, making fine-grained adjustments more challenging. Also, since low-frequency components typically require more energy [52], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details. As shown in fig.˜2(a), higher frequency components converge after their counterpart. Hence, in this work, we generate the entire basis map $b_{i}=M_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
(a) $b_{i}(\mathbf{x})=\cos\left[({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{i}}x+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v_{i}}y)-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\xi_{i}}\right]$
(b) $b_{i}(\mathbf{x})=\cos\left[(u_{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x}+v_{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}y})-\xi_{i}\right]$
(c) $b_{i}=\cos\left(\alpha_{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{i}}\right)+\cos\left(\alpha_{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{i}}\right)$
Figure 2: Comparison of different trainable basis settings We test on single-image regression with $f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,b_{i}(\mathbf{x})$ with equal parameter count. (a) fixed sinusoidal basis favoring low–frequency content, eq.˜3; (b) learnable (i.e. requires gradient) coordinate mapping with moderate gains; (c) Learned Non-uniform Basis $M_{i}\in\mathbb{R}^{T\times T}$ converges more synchronously. Note that $\alpha_{1},\alpha_{2}$ are the frequencies, functionally similar to $u,v$.
To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map, and combine them. In this work, the composition process happens in the end with a projection function $\mathcal{P}$ for RGB images $\hat{I}$:
$\small\hat{I}(\mathbf{x})=\mathcal{P}\Bigl(\textbf{Concat}_{i=1}^{N}\Bigl\{c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl\}\Bigl).$ (5)
@@ 4.1 Formulation of Factorized Features
Next, we discuss more improvements on frequency decomposition to model image details.
Refer to caption Refer to caption
(a) eq.˜5 (b) eq.˜6
Refer to caption Refer to caption
(c) eq.˜7 with only $\psi$ (d) eq.˜7 with both $\psi$ and $\alpha$
Figure 3: Visualization of different factorization methods This figure shows the difference between vanilla fields, coordinate transformation, and multi-frequency modulation. Specifically, (a) is a vanilla basis-coefficient field, and (b) adds sawtooth transformation. We can see in (b) that the coordinate transformation explicitly models a patch-like pattern, e.g., the second plot from the top is divided into two periods. By enforcing such frequency components, the models can decompose the signal effectively. Next, (d) is our full formulation, while (c) has no $\alpha$. Intuitively, from (b) to (c), applying $\psi$ (we use $\cos$ here) should not have much effect on the performance since we do not have any constraint on these learnable bases. With the introduction of $\alpha$, the models are forced to attend to signal components of different frequencies.
Coordinate Transformation. The success of the Fourier-based representation originates from its sinusoidal decomposition. By using such periodic functions, one can effectively analyze and reconstruct signals with the enforcement of high- and low-frequency components. However, the formulation in eq.˜5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in fig.˜3((a) vs (b)) we witness performance loss due to lack of such guidance. Inspired by position encoding [78, 11, 81, 7, 9] and Fourier feature mapping [91], we introduce coordinate transformation function $\gamma_{i}$ on $\mathbf{x}$ before sampling basis $b_{i}$:
$\small\hat{I}(x)=\mathcal{P}\Bigl(\textbf{Concat}_{i=1}^{N}\Bigl\{c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl\}\Bigl).$ (6)
In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$; visualization is also provided in the supplementary material. We argue that the function explicitly imposes patch-like periodic constraints and, therefore, enforces models to learn such repetitive patterns.
Multi-frequency Modulation. [85] has shown that neural networks are biased towards learning low-frequency content. Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in fig.˜3c and fig.˜3d:
$\small\hat{I}(x)=\mathcal{P}\Bigl({\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl\{c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl\}\Bigl),$ (7)
@@
Alternatively, eq.˜7 can be interpreted as making the spatial coordinates $x,y$ of $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ in eq.˜3 learnable instead of the frequency components $u,v$. To validate this, we conduct experiments (fig.˜2(b)) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
To this end, our full Factorized Features representation is formulated in eq.˜7, visualized in fig.˜4.
Figure 4: The Proposed Factorized Features. In this figure $\alpha=\{1,16\}$, $\psi=\text{sin}(\cdot)$, and that $\gamma_{i}(x)$ is coordinate transformation. We can see in this figure with sawtooth transformation $\gamma$ the feature is prompted to learn pattern from different scale in a patch-like manner, e.g. $\gamma_{2}$ makes the same basis repeat four times in the grid. However, only $\gamma$ is not enough for very fine details; therefore, we introduce $\alpha$ and $\psi$. Each basis is forced to accommodate both high-frequency (large $\alpha$ ) and low-frequency (small $\alpha$) components together with periodic functions $\psi$, i.e., larger $\alpha$ will make $\psi$ oscillate sharper.With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression.